# Timelike electromagnetic form factors of hyperons at large $q^{2}$

### v1

@@ Abstract
In the last few years there has been considerable progress in the study of the electromagnetic form factors of baryons in the timelike region, through electron-positron scattering, with increasing square transfer momentum $q^{2}$. The modulus of the electric ($G_{E}$) and magnetic ($G_{M}$) form factors has been measured for nucleons, hyperons and other baryons at BaBar, CLEO, Belle and BESIII. The novel measurements motivated the extension of a covariant quark model, developed to the spacelike region ($q^{2}\leq 0$), to the timelike region, without any further parameter fitting. The extension is based on asymptotic relations derived from analyticity and unitarity, valid for the large-$q^{2}$ region. We use the model to make predictions for the effective form factor $|G|$ (combination of $G_{E}$ and $G_{M}$) and the ratio $|G_{E}/G_{M}|$ for spin 1/2 hyperons at large $q^{2}$ (above 10 GeV2). Our calculations are in good agreement with the data from CLEO and BESIII for $\Lambda$, $\Sigma^{+}$ and $\Xi^{-}$ above $q^{2}=15$ GeV2. Upcoming data for $\Sigma^{0}$, $\Sigma^{-}$ and $\Xi^{-}$ at large $q^{2}$ may be used to further test our predictions. We also compare our model calculations with the scarce available data for $|G_{E}/G_{M}|$. We conclude that the present $q^{2}$ range is not large enough to test our calculations, but that a more definitive test can be made by experiments above $q^{2}=20$ GeV2.
@@ 1 Introduction and formalism
[…]
@@ 2 Numerical results
In the covariant spectator quark model the electromagnetic transition current is determined using relativistic impulse approximation for the photon interaction with a single quark, while integrating over the degrees of freedom of the spectator quark-pair [12, 13, 14, 16]. The quark internal structure associated with the gluon and quark-antiquark dressing is parametrized by constituent quark form factors. The quark-diquark radial wave functions are determined phenomenologicaly using lattice QCD data, based on the SU(3) flavor symmetry structure that is broken explicitly. In the low-$Q^{2}$ region, we consider also effective descriptions of the meson cloud dressing of the baryon states [13, 16].
@@ Acknowledgments
G.R. and M.K.C. were supported the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (Grant No. NRF–2021R1A6-A1A03043957). M.T.P was supported by the Portuguese Science Foundation FCT under project CERN/FIS-PAR/0023/2021, and FCT computing project 2021.09667. K.T. was supported by CNPq, Brazil, Processes No. 304199/2022-2, FAPESP Process No. 2023/07313-6, and by Instituto Nacional de Ciência e Tecnologia – Nuclear Physics and Applications (INCT-FNA), Brazil, Process No. 464898/2014-5.

### v2

@@ Abstract
In the last few years there has been considerable progress in the study of the electromagnetic form factors of baryons in the timelike region, through electron-positron scattering, with increasing squared transfer momentum $q^{2}$. The modulus of the electric ($G_{E}$) and magnetic ($G_{M}$) form factors has been measured for nucleons, hyperons and other baryons at BaBar, CLEO, Belle and BESIII. The novel measurements motivated the extension of a covariant quark model, developed to the spacelike region ($q^{2}\leq 0$), to the timelike region, without any further parameter fitting. The extension is based on asymptotic relations derived from analyticity and unitarity, valid for the large-$q^{2}$ region. We use the model to make predictions for the effective form factor $|G|$ (combination of $G_{E}$ and $G_{M}$) and the ratio $|G_{E}/G_{M}|$ for spin 1/2 hyperons at large $q^{2}$ (above 10 GeV2). Our calculations are in good agreement with the data from CLEO and BESIII for $\Lambda$, $\Sigma^{+}$ and $\Xi^{-}$ above $q^{2}=15$ GeV2. Upcoming data for $\Sigma^{0}$, $\Sigma^{-}$ and $\Xi^{-}$ at large $q^{2}$ may be used to further test our predictions. We also compare our model calculations with the scarce available data for $|G_{E}/G_{M}|$. We conclude that the present $q^{2}$ range is not large enough to test our calculations, but that a more definitive test can be made by experiments above $q^{2}=20$ GeV2.
@@ 1 Introduction and formalism
$\frac{G_{E}}{G_{M}}=\frac{|G_{E}|}{|G_{M}|}e^{i\Delta\Phi},\hskip 56.9055pt{\rm Re}\left(\frac{G_{E}}{G_{M}}\right)=\frac{|G_{E}|}{|G_{M}|}\cos\Delta\Phi.$ (2)
@@ 2 Numerical results
In the covariant spectator quark model the electromagnetic transition current is determined using relativistic impulse approximation for the photon interaction with a single quark, while integrating over the degrees of freedom of the spectator quark-pair [12, 13, 14, 16]. The quark internal structure associated with the gluon and quark-antiquark dressing is parametrized by constituent quark form factors. The quark-diquark radial wave functions are determined phenomenologically using lattice QCD data, based on the SU(3) flavor symmetry structure that is broken explicitly. In the low-$Q^{2}$ region, we consider also effective descriptions of the meson cloud dressing of the baryon states [13, 16].
@@ Acknowledgments
G.R. and M.-K.C. were supported by the National Research Foundation of Korea (Grant No. RS-2021-NR060129). M.T.P was supported by the Portuguese Science Foundation FCT under project CERN/FIS-PAR/0023/2021, and FCT computing project 2021.09667. K.T. was supported by CNPq, Brazil, Processes No. 304199/2022-2, FAPESP Process No. 2023/07313-6, and by Instituto Nacional de Ciência e Tecnologia – Nuclear Physics and Applications (INCT-FNA), Brazil, Process No. 464898/2014-5.